The sum of 4 times a number and 49 is the negative product of 3 and the number?
Answer by checkley77(12844)

4x+49=-3x
4x+3x=-49
7x=-49
x=-49/7
x=-7 answer.
Proof
4*-7+49=-3*-7
-28+49=21
21=21